# Boosting Generalization in Diffusion-Based Neural Combinatorial Solver via Energy-guided Sampling

### v1

@@ 6.1 Generalization Results
Learning-based AM (Greedy) 2.76% 0.02s 18.20% 0.07s 28.98% 0.15s 16.65% 3.5 days
@@
DIF-Trans (Ours) 4.91% 8.32s 8.77% 10.46s 18.02% 20.18s 10.56% 0 day
Table 1: Comprehensive evaluation of cross-scale generalization capabilities across different solver categories on PCTSP instances. Comparison includes exact solvers (Gurobi), OR-based heuristics (OR-Tools, ILS), learning-based models (AM, MDAM, ASP, AM-FT, DIF-Trans). Performance metrics include optimality gap, inference time, and training time. The results marked with * are reported from [10]. The proposed DIF-Trans achieves competitive performance while requiring no training.
@@
Learning-based AM (Greedy) 5.78% 0.04s 12.10% 0.09s 28.75% 0.19s 15.54% 3.2 days
@@
DIF-Trans (Ours) 3.59% 7.89s 6.83% 10.77s 14.33% 19.06s 8.25% 0 day
Table 2: Comprehensive evaluation of cross-scale generalization capabilities across different solver categories on OP instances. Comparison includes exact solvers (Gurobi), OR-based heuristics (Compass, Tsili), learning-based models (AM, AM-FT, DIF-Trans). The results marked with * are reported from [11].
Experimental results presented in Table 2 and Table 2 demonstrate the superior generalization capabilities of our approach. DIF-Trans achieves competitive average optimality gaps ($10.56\%$ for PCTSP and $8.25\%$ for OP) across all problem scales while maintaining zero-shot inference capability. It eliminates the substantial computational overhead (3-5 days for existing methods) and well-labeled training data associated with problem-specific training. While OR-based heuristics like Compass showcase strong performance only on specific problem types, DIF-Trans’s ability to leverage pre-trained diffusion models for different CO problem variants without additional training offers both scalability and adaptability in NCO solvers.
@@ 6.2 Ablation Studies
To validate the effectiveness of our energy-guided sampling framework in cross-problem transfer scenarios, we conduct comprehensive ablation studies while maintaining consistent problem scales. As demonstrated in Table 3, our method achieves significant zero-shot performance improvements when applied to the pre-trained DIFUSCO model on TSP. These improvements are evident across all PCTSP and OP problem scales, manifesting in both enhanced solution quality and reduced optimality gaps compared to traditional OR-based solvers.
Figure 3: Trade-off between optimality gap and inference time of energy-guided sampling with respect to rewriting iterations on PCTSP.
Figure 4: Optimality gap and inference time on OP.
The cross-problem transfer capabilities exhibit progressive enhancement through our iterative rewriting mechanism, where each round’s noisy data $\mathbf{x}_{T}$ is initialized using the previously obtained optimal solution. Figures 4 and 4 quantitatively demonstrate the solution adaptation trajectory as iterations proceed. While this iterative sampling paradigm causes additional computational overhead during inference, our training-free approach maintains computational efficiency and scalability compared to conventional OR solvers, which are inherently constrained by strict problem formulations and scale limitations.
@@
PCTSP-20 PCTSP-50 PCTSP-100
[…]
DIFUSCO 3.78 19.21% 1.04s 5.20 15.97% 1.35s 8.14 35.61% 2.85s
DIF-Trans (Ours) 3.32 4.83% 7.20s 4.69 3.79% 9.15s 6.67 12.31% 18.41s
OP-20 OP-50 OP-100
[…]
DIFUSCO 9.25 12.48% 1.51s 25.60 13.45% 1.88s 45.27 23.72% 2.97s
DIF-Trans (Ours) 10.19 3.59% 6.97s 28.15 4.82% 10.44s 53.35 9.85% 19.20s
Table 3: Zero-shot cross-problem transfer performance comparison between TSP-trained DIFUSCO and our DIF-Trans approach on PCTSP and OP instances. The results show solution cost, optimality gap, and inference time across three problem scales (20, 50, and 100 nodes).
Figure 5 demonstrates the effect of the temperature parameter in modulating the guidance strength of our framework. Higher temperature values effectively preserve the structural information in the original problem instances, whereas lower temperatures adapt to the characteristics of new problem instances. This temperature-dependent behavior enables fine-grained control over the balance between maintaining prior knowledge and new problem features.
7 Conclusions and Limitations
In this work, we introduced an energy-guided sampling framework enabling zero-shot cross-problem generalization for diffusion-based solvers. Through an energy-based guidance mechanism during inference, our approach successfully transfers pre-trained diffusion models to address TSP variants without requiring additional training. We conducted theoretical analysis in this scenario to deepen our understanding of the problem transfer capability. Extensive experiments on OP and PCTSP demonstrate that our framework achieves competitive performance against existing learning-based methods across various problem scales.
This work demonstrates the considerable potential of adapting diffusion-based generative neural solvers to address real-world combinatorial optimization problems, particularly those involving dynamic constraints and objectives. Our approach reduces the computational costs of retraining large-scale neural solvers by providing a flexible, off-the-shelf sampling scheme that enhances generalization capabilities. Nevertheless, several limitations and opportunities for future research consideration still exist. First, while our method exhibits promising results on TSP variants, its applicability to a broader spectrum of combinatorial optimization problems beyond routing domains remains to be validated. Second, the computational complexity introduced by the energy-guided sampling process during inference requires further optimization. Future research directions could focus on extending the framework to diverse combinatorial optimization paradigms and developing more efficient guided-sampling scheme.
Appendix A Appendix
A.1 Proof of Theorem 4.2
@@ Appendix A Appendix
Definition A.1 (Marginal Decrease).
@@
Assumption A.2 (PCTSP Setting).
@@
The above identical reward assumption can be relaxed by considering a node-weighted TSP, which can be transformed into a standard TSP with some effort. Then, we can apply the following arguments for OP to build a connection with the optimal solutions of standard TSP.
@@
A.2 Energy Formulations for TSP Variants
In our experimental setup, we employ an anisotropic graph neural network architecture enhanced with edge-gating mechanisms [37, 38]. This encoder transforms the input 2D coordinates and TSP variant paths into a binary edge-selection graph. Additional problem-specific information - such as prizes and penalties in PCTSP and prizes in OP - is incorporated into the gradient calculation of the guidance objective function. The constraint coefficients $\beta$ are chosen to be $1.1$ in experiments.
Proposition A.4 (Traveling Salesman Problem (TSP)).
@@ A.2 Energy Formulations for TSP Variants
$f(\mathbf{x};G)$ $=f_{\text{cost}}(\mathbf{x};G)+\beta\cdot f_{\text{valid}}(\mathbf{x};G)$
@@
Proposition A.5 (Prize Collecting TSP (PCTSP)).
$f(\mathbf{x},\mathbf{y};G)$ $=f_{\text{cost}}(\mathbf{x},\mathbf{y};G)+\beta\cdot f_{\text{valid}}(\mathbf{x},\mathbf{y};G)$
$f_{\text{valid}}(\mathbf{x},\mathbf{y};G)$ $=\max(0,R-\sum_{v\in V}r_{v}y_{v})$
Proposition A.6 (Orienteering Problem (OP)).
$f(\mathbf{x},\mathbf{y};G)$ $=f_{\text{cost}}(\mathbf{x},\mathbf{y};G)+\beta\cdot f_{\text{valid}}(\mathbf{x},\mathbf{y};G)$
$f_{\text{valid}}(\mathbf{x},\mathbf{y};G)$ $=\max(0,\sum_{e\in E}w_{e}x_{e}-B)$

### v2

@@ 5.1 Cross-problem and Cross-scale Generalization Results
Learning AM (Greedy) 2.76% 0.02s 18.20% 0.07s 28.98% 0.15s 16.65% 3.5 days
@@
DIFU-Ada (Ours) 4.45% 1.80s 8.43% 1.92s 16.22% 2.87s 9.70% 0 day
Table 2: Comprehensive evaluation of cross-scale generalization capabilities across different solver categories on PCTSP instances. Comparison includes exact solvers (Gurobi), OR-based heuristics (OR-Tools, ILS), learning-based models (AM, MDAM, ASP, AM-FT, DIFU-Ada). Performance metrics include optimality gap, inference time, and training time. The results marked with * are reported from [11]. The proposed DIFU-Ada achieves competitive performance while requiring no training.
@@
Learning AM (Greedy) 5.78% 0.04s 12.10% 0.09s 28.75% 0.19s 15.54% 3.2 days
@@
DIFU-Ada (Ours) 3.44% 2.01s 6.11% 2.29s 12.21% 2.85s 7.26% 0 day
Table 3: Comprehensive evaluation of cross-scale generalization capabilities across different solver categories on OP instances. Comparison includes exact solvers (Gurobi), OR-based heuristics (Compass, Tsili), learning-based models (AM, AM-FT, DIFU-Ada). The results marked with * are reported from [12].
Figure 4: Impact of the number of recursive travel steps in the inference time adaptation on the trade-off between optimality gap (%) and inference time (s) for PCTSP and OP.
@@ 5.2 Ablation Studies
The DIFU-Ada framework integrates two key components: energy-guided sampling and recursive denoising-decoding travel strategy. To assess the contribution of each, we conducted ablation studies. Figure 4 quantitatively illustrates the impact of the recursive denoising-decoding strategy, demonstrating a progressive enhancement in cross-problem transfer capabilities as the number of recursive iterations increases. Figure 5 presents a analysis of performance with or without energy-guided sampling during the recursive iterations. The performance gap observed in these studies show the efficacy of both components and highlight their distinct roles in the overall effectiveness to achieve zero-shot cross-problem transfer.
Moreover, Figure 6 shows the effect of the temperature parameter in controlling the guidance strength of adaptation. Higher temperature effectively preserve the structural information in the original problem instances, whereas lower temperatures adapt to the characteristics of new problem instances. This temperature-dependent behavior enables fine-grained control over the balance between maintaining prior knowledge and new problem features. In Appendix F, Figure 7, the results on varying hyperparameter $\mu$ in energy function show that the performances are not sensitive to the choice of $\mu$, thus we choose the optimal $\mu=1$ empirically for all problems.
Figure 5: The Optimality Gap of performing recursive travel strategy with or without energy-guided sampling on PCTSP and OP.
@@
6 Conclusions
In this work, we introduced an Inference Time Adaptation framework (DIFU-Ada) enabling zero-shot cross-problem transfer for diffusion-based solvers. Through an integration of energy-guided sampling and recursive denoising-decoding travel strategy, our approach successfully transfers pre-trained diffusion models to generate high-quality solutions for TSP variants without additional training. We also provided theoretical analysis in this scenario to deepen our understanding of the problem transfer capability in pre-trained diffusion-based NCO solvers. Numerical experiments on different scales of PCTSP and OP demonstrate that our framework achieves competitive zero-shot performance against existing learning-based methods across various scales.
This work demonstrates the considerable potential of adapting diffusion-based generative NCO solvers to address real-world combinatorial optimization problems, particularly those involving dynamic constraints and objectives. Our approach reduces the computational costs of retraining large-scale neural solvers by providing a flexible, off-the-shelf sampling scheme that enhances both cross-problem and cross-scale generalization capabilities. Nevertheless, several limitations and directions for future research still exist. While our method exhibits promising results on TSP variants, its applicability to a broader spectrum of combinatorial optimization problems beyond routing domains remains to be validated.
Appendix A Problem Descriptions of TSP Variants
Appendix B Proof of Theorem 4.4
@@ Appendix A Problem Descriptions of TSP Variants
Definition B.1 (Marginal Decrease).
@@
Assumption B.2 (PCTSP Setting).
@@
Note that the above assumptions can be relaxed by considering a node-weighted TSP, which can be transformed into a standard TSP as discussed in Appendix C.
@@
Appendix C Relaxing Assumptions B.2 and B.3
Our analysis in Theorem 4.4 can be extended to the cases with minimum prize constraints or non-uniform rewards by considering node-weighted TSP formulations. The key is that the optimal solutions of PCTSP/OP are the optimal solutions of certain node-weighted TSP over a subgraph as shown below.
(i) For the general case of OP, we can reformulate the problem by assigning each node $i$ a score $s_{i}$ equal to its negative reward and setting all edge costs to zero. Meanwhile, we maintain the edge cost function $w_{e}$ solely to represent the budget constraint of OP. This transforms OP into a node-weighted TSP with an additional budget constraint.
(ii) Similarly, for PCTSP, we assign $s_{i}=-p_{i}$ (the negative penalty of node $i$) while preserving the original edge costs. The optimal solution to PCTSP then corresponds to that of a node-weighted TSP with minimum prize constraints.
Since the pre-trained model was trained on standard TSP instances, we can convert node-weighted TSP to a standard one by splitting each node $i$ into two nodes $i_{in}$ and $i_{out}$, and adding a directed edge from $i_{in}$ to $i_{out}$ with weight $s_{i}$. Given this conversion, we could explain the effectiveness of our method as follows: the energy-guided sampling implicitly considers the node-weighted TSP and guides the generation to zoom in a subgraph of the node-weighted graph which respects the additional constraints.
Appendix D Adaptation to TSP with Time Windows (TSP-TW)
TSP with Time Windows (TSP-TW) is a variant of TSP that requires each node $i$ to be visited within a given time range $[e_{i},l_{i}]$.
Our approach can also handle TSP-TW through a time-expanded graph transformation:
For each node $i$ with time window $[e_{i},l_{i}]$, we create multiple time-indexed replicas $(i_{e_{i}},i_{e_{i}+1},...,i_{l_{i}})$. The original graph’s edges are then transferred to this expanded representation according to temporal feasibility - if node $i$ connects to node $j$ in the original graph, we connect $i_{t}$ to $j_{t+1}$ in the expanded graph (assuming unit travel time for simplicity).
This transformation converts the time window constraints into a structural pattern within the expanded graph. The TSP-TW solution corresponds to a standard TSP tour on this expanded graph with one critical constraint: exactly one replica of each original node must be visited, expressed as:
$\sum_{t\in[e_{i},l_{i}]}\sum_{(j,i_{t})\in E}x_{j,i_{t}}=1.$
This constraint can be directly encoded as an energy function in our framework. Our method then proceeds by:
  • Sampling initial TSP solutions on the time-expanded graph
  • Using energy-guided sampling to enforce the single-visit-per-node constraint
  • Converging toward feasible TSP-TW solutions
Through this transformation, TSP-TW effectively becomes a "subgraph TSP" scenario similar to that of PCTSP/OP, demonstrating the flexibility of our energy-guided approach for handling complex constraints without requiring problem-specific training.
Appendix E Energy Formulations for TSP
In our experimental setup, we employ an anisotropic graph neural network architecture enhanced with edge-gating mechanisms [45, 46]. This encoder transforms the input 2D coordinates and TSP variant paths into a binary edge-selection graph. Additional problem-specific information - such as prizes and penalties in PCTSP and prizes in OP - is incorporated into the gradient calculation of the guidance objective function. The constraint coefficients $\mu$ are chosen to be $1$ in experiments.
Proposition E.1 (Traveling Salesman Problem (TSP)).
@@ Appendix E Energy Formulations for TSP
$\phi(\mathbf{x};G)$ $=f_{\text{cost}}(\mathbf{x};G)-\mu\cdot\log(-g(\mathbf{x};G))$
@@
Appendix F Additional Experimental Results
PCTSP-500 PCTSP-1000
[…]
OR Tools* 15.0 1h 24.9 1h
OR Tools (more iterations)* 14.4 16h 20.6 16h
AM* 19.3 14m 34.8 23m
MDAM* 14.8 2.8m 22.2 17m
GLOP-S* 14.3 1.5m 19.8 2.5m
DIFU-Ada (Ours) 14.9 7.2m 23.5 9.9m
Table 4: Comparison of various methods on 128 instances of PCTSP-500 and PCTSP-1000. The data generation and evaluations are conducted following Ye et al. (2024). Results marked with * are taken directly from Ye et al. (2024).
Training Method Testing Dataset
TSP-20 TSP-50 TSP-100
TSP-20 DIFUSCO 3.14, 0.23% 5.75, 0.43% 7.90, 1.81%
T2T 3.14, 0.12% 5.71, 0.19% 7.82, 0.87%
Fast T2T 3.12, 0.03% 5.69, 0.00% 7.76, 0.08%
Ours 3.14, 0.15% 5.71, 0.22% 7.83, 0.95%
TSP-50 DIFUSCO 3.14, 0.25% 5.69, 0.09% 7.87, 1.44%
T2T 3.14, 0.11% 5.69, 0.02% 7.80, 0.55%
Fast T2T 3.12, 0.05% 5.69, 0.00% 7.76, 0.08%
Ours 3.14, 0.14% 5.69, 0.05% 7.81, 0.60%
TSP-100 DIFUSCO 3.16, 0.44% 5.70, 0.25% 7.78, 0.23%
T2T 3.14, 0.21% 5.70, 0.11% 7.77, 0.08%
Fast T2T 3.12, 0.10% 5.69, 0.01% 7.76, 0.01%
Ours 3.14, 0.20% 5.71, 0.25% 7.77, 0.09%
Table 5: Results of various diffusion-based methods [6, 7, 8] on different TSP testing datasets (TSP-20, TSP-50, TSP-100), only considering cross-scale generalization capability.
Figure 7: Impact of the energy function hyperparameter $\mu$ on the optimality gap for PCTSP-50 and OP-50.